# Fundamental physical constants set the observability and operation of phase and other transitions and increase entropy

### v4

@@ 3.5 Metastable states and chemical reactions
In the previous three sections, we discussed the observability of phase transitions and the existence of different states as a result of these phase transitions. The usual statistical-mechanical treatment of these transitions assumes equilibrium phases corresponding to the global minimum of the free energy. However, many simple molecular compounds of light elements (e.g., most hydrocarbons, nitrogen oxides, hydrates, carbides, carbon oxide, alcohols, glycerin) are metastable and do not correspond to the lowest minimum of the free energy. Nevertheless, long-lived metastable phases in a certain range of pressure and temperature undergo reversible phase transitions which are consistent with laws of equilibrium thermodynamics: the phase transition corresponds to the free energy crossover (see Ref. [52] for review). When discussing phase transitions in these compounds, the tacit assumption is the one we discussed in Section 3.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
@@
Transitions between metastable states involve the same atomistic processes as transitions between equilibrium states and energies related to the chemical bond energy set by FPCs. As in the equilibrium case, these transitions correspond to the free energy crossover. Hence, similarly to what we saw for equilibrium transitions in earlier sections, the observability and operation of transitions between different metastable states and hence the existence of numerous metastable phases is set by FPCs.
Similarly to what we saw in Section 3.2, the entropy increases due to new metastable states emerging as a result of these transitions because the states formed at higher temperature tend to be more symmetric with more open structure and have larger phase space as a result (this larger phase space and higher entropy become frozen in the low-temperature metastable phase).
The same assertion of observability of transitions and different phases applies to chemical reactions and reactions products. Chemical reactions involve changes of chemical bonding whose characteristic energies are set by FPCs. Hence these reactions are governed by the interplay of characteristic cohesive energies set by FPCs and environmental parameters. For example, much larger cohesive energy due to different FPCs would imply that chemical reactions and their products (including those essential to life forms) would be unobservable at given environmental parameters.
@@ 4 Fundamental constants and entropy increase
We consider that understanding the current values of FPCs requires the discussion of their variation. Theories and observations of this variation were previously discussed [54, 55, 56, 57, 58, 59, 60, 6, 13, 61]. Anthropic arguments use this variation to rationalise the observed values of FPCs [9, 10, 13]. A different perspective should be noted: a final theory of most fundamental laws of nature, when found, would be rigid (“logically isolated”) [33] and guided by a unique design and an exquisite symmetry [62]. FPCs would then be calculable from this theory.
@@
In Section 3, we discussed different cases where FPCs, by enabling phase transitions and different phases at given environmental parameters, increase the entropy. We saw that the current values of FPCs enable phase and other transitions, giving rise to different states and the associated increase of the phase space and entropy. We also saw that different values of FPCs would disable these transitions and different phases, resulting in smaller entropy. This suggests that the current values of FPCs give rise to entropy increase.
@@
We discussed the entropy increase due to FPCs in condensed matter systems, however this increase is a general effect and can be applied to different length and energy scales in the Universe on general grounds. This is supported by a more specifc observation. The temperature of transitions and reactions at different energy scales is related to the energy of interaction between constituents and the relevant cohesive energy at each scale. These energies depend on FPCs, implying that FPCs enable transitions and reactions at those scales as they do in condensed matter. Similarly to condensed matter systems, we expect the new states emerging as a result of these transitions to have higher entropy. Hence we expect that FPCs act to increase the entropy at different levels of structure in the Universe in Figure 1.
@@
The assertion of the entropy increase includes the usual proviso that we consider a subsystem or a collection of subsystems in the Universe which are closed or conditionally closed [36]. We also note that the environmental parameters can be variable in some processes. For example, nuclear synthesis reactions in a forming star depend on (environmental) pressure and temperature, which, in turn, depend on the mass of collapsing interstellar gas. This variability can affect transitions and reactions and the properties of emergent structures. In some processes such as the above example of nuclear synthesis in stars, the environmental parameters can also change as a result of reactions. Nevertheless, the transitions and reactions are enabled or disabled by FPCs at varying environmental parameters as they are at fixed parameters.
@@
We recall that the discussion in Section 3 included long-lived metastable states. We saw that at given environmental parameters, the collection of these states increase the phase space and entropy. As recently reviewed [63], metastable states are ubiquitous at different scales in the Universe including nuclei and astronomical objects (energy barriers setting the life time of these states can vary in a wide range 10-7 to 109 eV per particle, with corresponding relaxation times often exceeding the age of the Universe [63]). These metastable states, similarly to those discussed in Section 3, emerge as a result of different transitions and reactions which are sensitive to the values of FPCs [6, 13] and can co-exist at the same environmental parameters.
The entropy increase enabled by FPCs can be discussed in terms of an evolution-like mechanism where entropy increase is analogous to selection due to environmental pressure in biological evolution. The mechanism of variability or fluctuations of FPCs, analogous to genetic mutation in biological evolution, and its operation are unknown, although several proposals were discussed (see, e.g. Refs. [64, 65, 66, 67, 68] and Ref. [61] for review).
@@
Cyclic or oscillatory models of the Universe (see, e.g., Refs. [69, 64, 65, 66, 67, 68, 70] and Ref. [71] for a review including epistemic and philosophical aspects) can account for inheriting FPCs from one cycle to the next and maintain fluctuations around their values in the previous cycle. In these models, the change of FPCs is often related to the smallest size of the Universe and very high energy including the Planck energy scale where unknown physics operates. The size of the oscillating Universe in Ref. [70] at the inversion point is set by the system energy as a parameter in the model and does not need to be small.
@@
We might ask whether the current FPCs give structures in the Universe that correspond to the maximised phase space and entropy or close to it. Adams concludes [13] that our Universe is not the most abundant and complex possible: somewhat different FPCs would give more complex and abundant structures (e.g. more baryons and stars, more efficient nuclear synthesis and so on). Another interesting observation in Ref. [13] is that our Universe does not lie at the centre of allowed range of FPCs. The evolutionary mechanism with the entropy increase as a selection principle has several potential implications for these observations. We first note that this selection applies to FPCs and FPC-enabled structures existing at the time (as is the case in biological evolution) rather than to currently observed structures. Second, the observed structures depend on the magnitude of fluctuations of FPCs: large fluctuations destroy the intricate network of constraints on the values of FPCs needed for different structures to exist [13], resulting in scarcity of structures and smaller phase space. Smaller fluctuations maintain the continuity of structures in different cycles. As a result, different structures can emerge along a sequence of cycles, with the entropy of these structures fluctuating along this sequence. Third, maximising a property (entropy) as a function of FPCs is subject to many constraints on their values. This invites an input from the optimisation theory.
@@
where $M_{BH}$ is the BH mass. However, small $\hbar$ and $c$ give large $\alpha$ in Eq. (2), breaking many constraints required for key processes of structure formation and disrupting these processes [6, 13]. This includes processes preceding the formation of BHs such as stellar evolution which affect BHs at later times [6, 13] (formation of large BHs is currently under discussion and includes the accretion involving smaller BHs). We also note that considering variation of the BH entropy due to different FPCs is unimportant for the evolution of FPCs at earlier stages when BHs may not have existed yet: as mentioned earlier, this evolution applies to structures existing at the time only.
This discussion and the conjecture of FPCs acting to increase the entropy need further study. This includes detailing entropy changes in processes at different levels of structure organisation in the Universe [6, 13], identifying sets of subsystems which can be considered closed or conditionally closed and where the entropy argument applies, clarifying the role of gravity [36] and so on. Given the scale of the problem we face in understanding the origin of fundamental constants [6, 15, 14], making new observations and asking new questions is useful for progress.
@@ 5 Discussion and summary
Here, we discussed a wider role played by FPCs in condensed matter physics: at given environmental parameters, FPCs define observability and operation of entire physical effects and phenomena. We gave examples of operation of first- and second-order phase transitions and emerging new phases, existence of numerous metastable phases, chemical reactions, their products and so on. We saw that were FPCs to take different values, transitions would be disabled at our environmental conditions and new different phases would not emerge. This suggests that the values of FPCs, by enabling transitions and new states emerging as a result of these transitions, act to increase the phase space and entropy.
This list of effects and phenomena in condensed matter can be probably extended significantly. The reason it is important to study how FPCs affect the operation and observability of different effects and phenomena in condensed matter is that this field is developed at the great level of detail. Physical models of condensed matter have been tested and cross-checked in numerous different experiments and simulations. In this sense, the insights into FPCs from condensed matter physics face less uncertainty and challenges as compared to models in high-energy physics and astrophysics (see Ref. [13] for a review of these challenges).
If the insights into the role of FPCs are found to be general enough, these insights can be applied to structures at other levels in the Universe. The insight that FPCs promote entropy increase is one such example. This and similar insights can provide useful links between different areas of physics and open up new discussions.
Last but not least, deeper understanding the role of FPCs in condensed matter physics is important on its own. A fairly recent realisation that FPCs put lower or upper bounds on key condensed matter properties has deepened our understanding of how these properties operate and benefited our physical models. These properties include viscosity [17], diffusion and spin dynamics [18, 19, 20, 21, 22], electron and heat transport [23, 24, 25, 26], thermal conductivity in insulators [27] and conductors [28], the speed of sound [29], elastic moduli including those in lower-dimensional systems [28, 30] and superconductivity [32]. Understanding how FPCs govern the observability and operation of entirely new effects and phenomena is another useful insight enhancing our understanding of condensed matter physics.
I am grateful to F. C. Adams, V. V. Brazhkin, M. T. Dove, M. Hutcheon, S. A. Kauffman, B. Monserrat, A. J. Phillips and C. J. Pickard for discussions and EPSRC for support.

### v5

@@ 3.5 Metastable states and chemical reactions
In the previous sections, we discussed phase transitions and the existence of different states as a result of these transitions. The usual statistical-mechanical treatment of these transitions assumes equilibrium phases corresponding to the global minimum of the free energy. However, many simple molecular compounds of light elements (e.g., most hydrocarbons, nitrogen oxides, hydrates, carbides, carbon oxide, alcohols, glycerin) are metastable and do not correspond to the lowest free energy minimum. Nevertheless, long-lived metastable phases in a certain range of pressure and temperature undergo reversible phase transitions which are consistent with equilibrium thermodynamics: the phase transition corresponds to the free energy crossover (see Ref. [63] for review). When discussing transitions in these systems, the tacit assumption is the one we mentioned in Section 3.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
@@
Transitions between metastable states involve the same atomistic processes as transitions between equilibrium states and energies related to the chemical bond energy set by FPCs. Hence, similarly to equilibrium transitions, the observability and operation of transitions between different metastable states and the existence of numerous metastable phases is set by FPCs. The entropy of new high-temperature metastable states emerging in these transitions is expected to be higher because these states tend to be more symmetric with more open structure, as is the case for equilibrium transitons.
The same assertion of observability of transitions and different phases applies to chemical reactions and their products. Chemical reactions involve changes of chemical bonds with characteristic energies set by FPCs. Hence these reactions are governed by the cohesive energies set by FPCs and environmental parameters. For example, much larger cohesive energy due to different FPCs would imply that chemical reactions and their products (including those essential to life forms) would be inoperable at given environmental parameters.
@@ 4 Fundamental constants and entropy increase
We consider that understanding the current values of FPCs requires the discussion of their variation. Theories and observations of this variation were previously discussed [35, 36, 37, 38, 39, 40, 41, 42, 6, 13, 4, 33]. Anthropic arguments use this variation to rationalise the observed values of FPCs [9, 10, 13]. A different perspective should be noted: a final theory of most fundamental laws of nature, when found, would be rigid (“logically isolated”) [43] and guided by a unique design and an exquisite symmetry [65]. FPCs would then be calculable from this theory.
@@
In Section 3, we discussed different cases where FPCs, by enabling phase transitions and different phases, increase the entropy. We saw that the current values of FPCs enable phase and other transitions, giving rise to different states and the associated increase of the phase space and entropy. We also saw that different values of FPCs would disable these transitions and different phases, resulting in smaller entropy. This suggests that the current values of FPCs give rise to entropy increase.
@@
We discussed the entropy increase in condensed matter systems, however this increase applies to different length and energy scales in the Universe on general grounds. Recall Eq. (16) giving the temperature of phase transitions as a function of FPCs in condensed matter. However, the temperature of transitions and reactions depend on FPCs at other energy and length scales as well [6, 13], enabling transitions and reactions at those scales. Similarly to condensed matter systems, we expect the new states emerging as a result of these transitions to have higher entropy. Hence we expect that FPCs promote entropy increase at different levels of structure in the Universe.
This entropy increase is specifically due to transitions and reactions and comes in addition to the conventional general mechanism of entropy increase (including in the second law of thermodynamics). In the conventional mechanism, $S$ of a closed system at a given energy $E_{0}$ is maximised at equilibrium because (a) the statistical probability $d\rho$ is
$d\rho={\rm const}\cdot\delta(E-E_{0})e^{S}\prod\limits_{i}dE_{i}$ (22)
where $E_{i}$ are energies of subsystems and (b) $d\rho$ is maximised when $E_{i}$ of all subsystems in their own equilibrium states are equal to their most probable equilibrium values $\bar{E_{i}}$, corresponding to full equilibrium [46]. As discussed in Section 3.2, the subsystem where transitions and reactions take place (phase-changing subsystem, PCS) is not closed and not in equilibrium because these transitions are enabled by the heat flow from another heat-supplying subsystem. However, subsystems can be identified which are (a) larger, include both subsystems and are closed and (b) in equilibrium at long post-reaction times [46]. The conventional mechanism of increase of $S$ and $d\rho$ applies to these closed subsystems, and our entropy increase in the PCS enabled by FPCs additionally increases the statistical probability in Eq. (22). As discussed below, identifying collections of subsystems which can be considered as conditionally closed at different levels of the Universe can be of interest in future work.
@@
We note that the environmental parameters can be variable in some processes. For example, nuclear synthesis reactions in a forming star depend on (environmental) pressure and temperature, which, in turn, depend on the mass of collapsing interstellar gas. This variability can affect transitions and reactions and the properties of emergent structures. In some processes such as the above example of nuclear synthesis in stars, the environmental parameters can also change as a result of reactions themselves. Nevertheless, the transitions and reactions are enabled or disabled by FPCs at varying environmental parameters as they are at fixed parameters.
@@
We recall the discussion of long-lived metastable states with higher entropy in Section 3. As recently reviewed [66], metastable states are ubiquitous at different levels of matter organization in the Universe including nuclei and astronomical objects. The energy barriers setting the lifetime of these states vary in a wide range 10-7 to 109 eV per particle, with relaxation times often exceeding the age of the Universe [66]. These metastable states, similarly to those discussed in Section 3, emerge as a result of different transitions and reactions which depend on the values of FPCs [6, 13] and can co-exist at the same environmental parameters.
The entropy increase enabled by FPCs can be discussed in terms of an evolution-like mechanism where this increase is analogous to selection in biological evolution due to environmental pressure. Based on this entropy increase and the associated increase of statistical probability (22), we conjecture that entropy increase is a selection principle for FPCs considered to vary in earlier discussions (see, e.g., Refs. [35, 36, 37, 38, 39, 40, 41, 42, 6, 13, 4, 33] and references therein). In other words, FPCs tend to give the most probable state. The mechanism of variability or fluctuations of FPCs, analogous to genetic mutation in biological evolution, is unknown, although several proposals were discussed [35, 36, 37, 38, 39, 40, 41, 42, 6, 13, 33]. The above conjecture is nevertheless interesting to consider because, if correct, it is analogous to introducing a Darwinian concept of a selection principle, the notion that brought new understanding even though the molecular DNA-based mutation mechanism was not known at the time and came about a century later only.
@@
Cyclic or oscillatory models of the Universe (see, e.g., Refs. [67, 35, 68, 69, 70, 71, 72] and Ref. [73] for a review including epistemic and philosophical aspects) can account for inheriting FPCs from one cycle to the next and maintain fluctuations around their values in the previous cycle. In these models, the change of FPCs is often related to the smallest size of the Universe and very high energy including the Planck energy scale where unknown physics operates. The size of the oscillating Universe at the inversion point in Ref. [72] is set by the system energy as a parameter in the model and does not need to be small.
@@
We might ask whether the current FPCs give structures in the Universe that correspond to the maximised phase space and entropy or close to it. Adams concludes [13] that the observable Universe is not the most abundant and complex possible: somewhat different values of FPCs would give more complex and abundant structures (e.g. more baryons and stars, more efficient nuclear synthesis and so on). Another interesting observation in Ref. [13] is that our Universe does not lie at the centre of allowed range of values of FPCs. The evolutionary mechanism with the entropy increase as a selection principle has several potential implications for these observations. We first note that this selection applies to FPCs and FPC-enabled structures existing at the time (as is the case in biological evolution) rather than to currently observed structures. Second, the observed structures depend on the magnitude of fluctuations of FPCs: large fluctuations destroy the intricate network of constraints on the values of FPCs needed for different structures to exist [13], resulting in scarcity of structures and smaller phase space. On the other hand, smaller fluctuations maintain the continuity of structures in different cycles. As a result, different structures can emerge along a sequence of cycles, with the entropy of these structures fluctuating along the sequence. Third, it follows that maximising entropy as a function of FPCs is subject to many constraints on their values. This invites an input from the optimisation theory.
@@
where $M_{BH}$ is the BH mass. However, small $\hbar$ and $c$ give large $\alpha$ in Eq. (2), breaking many constraints required for key processes of structure formation and disrupting those processes [6, 13]. This includes processes preceding the formation of BHs such as stellar evolution which affect BHs at later times [6, 13] (formation of large BHs is currently under discussion and includes the accretion involving smaller BHs). We also note that considering variation of the BH entropy due to different FPCs is unimportant for the evolution of FPCs at earlier stages when BHs may not have existed yet: as mentioned earlier, this evolution applies to structures existing at the time only.
The discussion of FPCs acting to increase the entropy and the conjecture of entropy increase as a selection principle for FPCs need further study. This includes detailing entropy changes in processes at different levels of structure organisation in the Universe [6, 13], identifying sets of subsystems which can be considered closed or conditionally closed and where the entropy argument applies, clarifying the role of gravity [46] and so on. Given the scale of the problem we face in understanding the origin of fundamental constants [6, 15, 14], making new observations and asking new questions is useful for progress.
@@ 5 Discussion and summary
Here, we discussed a much wider role played by FPCs in condensed matter physics: at given environmental parameters, FPCs define observability and operation of entire physical effects and phenomena. Our examples included the operation of different phase transitions and emerging new phases, existence of different states of matter, metastable phases, chemical reactions, their products and so on. These results were based on a theory of each physical effect and process rather than on a dimensional analysis as might have been construed. Ref. [31] (pages 479-480) lists important differences between such a theory-based approach and a dimensional analysis.
We saw that were FPCs to take different values, transitions would be disabled at our environmental conditions and new different phases would not emerge. This suggests that the values of FPCs, by enabling transitions and new states emerging as a result of these transitions, act to increase the phase space and entropy.
The list of effects and phenomena in condensed matter physics discussed here can be probably extended significantly. The reason it is important to study how FPCs affect the operation and observability of different effects and phenomena in condensed matter is that this field is developed at the great level of detail. Physical models of condensed matter have been tested and cross-checked in numerous different experiments and simulations. In this sense, the insights into FPCs from condensed matter physics face less uncertainty and challenges as compared to models in high-energy physics and astrophysics (see Ref. [13] for a review of these challenges).
If the insights into the role of FPCs in condensed matter physics are found to be general enough, these insights can be applied to structures at other levels in the Universe. An example is the insight that FPCs promote entropy increase, or the related conjecture of the selection principle for FPCs. These and similar insights can provide useful links between different areas of physics and open up new discussions.
Last but not least, deeper understanding the role of FPCs in condensed matter physics is important on its own. A fairly recent realisation that FPCs put lower or upper bounds on key condensed matter properties has deepened our understanding of how these properties operate and benefited our physical models. These properties include viscosity [17], diffusion and spin dynamics [18, 19, 20, 21, 22], electron and heat transport [23, 24, 25, 26, 27], thermal conductivity in insulators [28] and conductors [29], the speed of sound [30], elastic moduli [29, 31], flow in cells [32, 33], vibration frequency in solids and superconducting temperature [34]. Understanding how FPCs govern the observability and operation of entirely new effects and phenomena is another useful insight enhancing our understanding of condensed matter physics.
I am grateful to F. C. Adams, V. V. Brazhkin, M. T. Dove, M. Hutcheon, S. A. Kauffman, Z.-K. Liu, B. Monserrat, A. J. Phillips and C. J. Pickard for discussions and EPSRC for support.